OpenStudy (anonymous):

What are the solutions of the system? y=x^2+4.6x+4.7 y=3.3x-3.7 A. (–3.47, 3.48) and (2.17, –15.17) B. (–3.47, –15.17) and (–2.17, –10.88) C. (–3.47, –15.17) and (2.17, 3.48) D. no solution A, B, C, or D? Please explain :)

OpenStudy (kira_yamato):

Equate the RHS of both together. x^2 + 4.6x + 4.7 = 3.3x - 3.7 x^2 + 1.3x - 1 = 0 Using quadratic formula \[x = \frac{-1.3 ± \sqrt{(1.3)^2 - 4(1)(8.4)}}{2(1)} = \frac{-1.3 ± \sqrt{-31.91}}{2}\] No soln

OpenStudy (anonymous):

Are you positive that there is no solution?

OpenStudy (anonymous):

Well then, thank you for all of your help.

OpenStudy (kira_yamato):

Because apparently there's no real solution... Because you cannot get a real no. by sqrt -ve no.